# Enhancing End Stage Renal Disease Outcome Prediction: A Multi-Sourced Data-Driven Approach

### v2

@@ 3.2.3 Additional Analyses
We analyzed instances where the model incorrectly predicted ESRD progression (type 1 error) and where it failed to predict ESRD progression in patients who eventually developed the condition (type 2 error). By comparing these error types, we sought to identify common factors or patterns contributing to misclassification, thus highlighting the model’s limitations and suggesting areas for refinement.
@@
Impact of Updated eGFR Equations on Racial Bias in CKD Predictions
The evolution of eGFR equations has aimed to address inherent limitations, particularly concerning racial bias. In 2020, the use of race-based components in eGFR calculations was scrutinized due to a lack of evidence of biological basis and the risk of perpetuating racial stereotypes and disparities in healthcare. In response, a 2021 workgroup led by the National Kidney Foundation and American Society of Nephrology recommended an updated CKD Epidemiology Collaboration (CKD-EPI) creatinine equation 21, which excluded the race coefficient to improve accuracy and equity 22. This adjustment aimed to enhance prediction precision, especially for minority populations 23; 24; 21.
Our analysis, using data from 2009 to 2018, initially applied the older eGFR equation. To assess the impact of the updated equation, we compared predictions between the old and new equations, focusing on improvements in accuracy and equity, particularly for minority groups.
@@ 4 Results
4.1 Data Integration and Cohort for Analysis
(a) Comparison of CKD stage distribution across claims, clinical, and merged datasets with a 24-month observation window.
@@ 4.1 Data Integration and Cohort for Analysis
(b) Variation in CKD stage distribution as a function of different observation window lengths.
@@
Figure 2: Overview of CKD Stage Distribution in Merged Data Across Observation Window Lengths.
@@
To maximize the utility of multi-sourced integrated data, we merged claims and clinical datasets using unique patient identifiers. The distribution of data for individual and merged datasets is presented in Fig. 2. The analysis cohort consists of patients identified at CKD stage 3, marking the beginning of data collection for each patient. As shown in Fig. 2(a), merging the datasets resulted in a significant reduction in the number of patients—from 5,518 and 6,842 in the claims and clinical datasets, respectively, to 1,422 in the merged dataset with a 24-month observation window. This reduction is mainly due to excluding patients who only had either claims or clinical data, resulting in a smaller but more comprehensive dataset for analysis. Fig. 2(b) illustrates the changes in the total number of patients and their distribution across CKD stages with varying observation window lengths. As the observation window extends from 6 to 30 months, there is a gradual decline in the number of patients available for analysis. This decline reflects the exclusion of patients who either lacked sufficient data within the specified window or progressed to ESRD during the period. Additionally, some patient losses may be due to mortality, but we lack the data to confirm this.
4.2 Performance Comparison of Models Across Datasets and Scenarios
4.2.1 Scenario 1: Model Robustness
This scenario focuses on evaluating the robustness and reliability of various models by training and testing them on a consistent dataset. We varied the training data length from 6 to 30 months within a fixed cohort to assess the models’ performance under different data availability conditions. This approach enabled us to evaluate the models’ stability and effectiveness across various observation windows, ensuring consistency and fair comparability.
@@ 4.2.1 Scenario 1: Model Robustness
Table 2: Comparison of AUROC and F1 scores across different models, using claims data-only, clinical data-only, and merged data in Scenario 1 (Model Robustness). This evaluation focuses on the stability and effectiveness of models across varying observation window lengths. Bold font is used to mark the best performance within each category (ML or DL methods), while bold and italic font highlights the best overall performance across all methods (ML + DL).

### v3

@@ 4.3.1 Analysis of Model Misclassifications
To evaluate the performance and robustness of our optimal model, we conducted an in-depth analysis of misclassifications by examining predictive probability distributions across the entire sample set. This analysis focused on type 1 (false positive) and type 2 (false negative) errors to understand discrepancies between predicted and actual outcomes. We analyzed instances where the model incorrectly predicted ESRD progression (type 1 error) and where it failed to predict ESRD progression in patients who eventually developed the condition (type 2 error). By comparing these error types, we sought to identify common factors or patterns contributing to misclassification, thus highlighting the model’s limitations and suggesting areas for refinement.
@@
4.3.2 Impact of Updated eGFR Equations on Racial Bias in CKD Predictions
The evolution of eGFR equations has aimed to address inherent limitations, particularly concerning racial bias. In 2020, the use of race-based components in eGFR calculations was scrutinized due to a lack of evidence of its biological basis and the risk of perpetuating racial stereotypes and disparities in healthcare. In response, a 2021 workgroup led by the National Kidney Foundation and American Society of Nephrology recommended an updated CKD Epidemiology Collaboration (CKD-EPI) creatinine equation 36, which excluded the race coefficient to improve accuracy and equity 37. This adjustment aimed to enhance prediction precision, especially for minority populations 38; 39.
Our analysis, using data from 2009 to 2018, applied both eGFR equations to the data. To assess the impact of the updated equation, we compared predictions between the old and new equations, focusing on improvements in accuracy and equity, particularly for minority groups.
It is important to note that adopting the updated equation may alter the cohort used for analysis, as the new eGFR values directly influence CKD stage classification. This could impact the cohort identification process, and lead to the reclassification of patients, particularly those near stage boundaries. Such changes may affect both the composition of the cohort and the model’s predictions.
@@ 5 Results
5.1 Cohort for Analysis & Data Characteristics
@@
(a) Variation in CKD stage distribution as a function of different observation window lengths.
@@
(b) Comparison of CKD stage distribution across claims, clinical, and merged data with a 30-month observation window.
Figure 3: Overview of CKD Stage Distribution in Merged Data Across Observation Windows.
@@
In this subsection, we present the trends observed during our data integration process and provide a detailed overview of the data characteristics for the final cohort utilized in our model performance analysis. Fig.3(a) illustrates the changes in patient distribution across CKD stages over varying observation window lengths. As the observation window extends from 6 to 30 months, there is a gradual decline in the number of patients available for analysis. This decline reflects the exclusion of patients who either lacked sufficient data within the specified window or progressed to ESRD during the period. Additionally, some patient losses may be due to mortality, but we lack the data to confirm this.
To ensure model consistency and fair comparability, we trained and tested on the same cohort. Recognizing the challenge of shrinking cohort size with extended observation windows, we focused on patients with at least 30 months of data following CKD stage 3 diagnosis and split the test data from this group. For the training data, we used the same cohort but with varying data lengths corresponding to the selected observation windows. This approach allowed us to explore the impact of different observation windows and identify the optimal one based on the performance on test data.
After establishing the cohort with at least 30 months of post-CKD stage 3 data, we compared the datasets derived from claims data only (5,518 patients), clinical data only (6,842 patients), and the merged dataset used for analysis. As shown in Fig. 3(b), merging the datasets resulted in a substantial reduction in the number of patients—to 1,422 patients with a 24-month observation window (which proved optimal in subsequent analyses) and 1,355 patients with a 30-month observation window. This reduction is mainly due to excluding patients who only had either claims or clinical data, resulting in a smaller but more comprehensive dataset for analysis.
Table 1: Data characteristics of the patient cohort for analysis
Characteristics Missing data (%) Progressed to ESRD (n = 86) Non-progressed to ESRD P-value
(n = 1,336)
Demographic
Age (years) 0 69.13 ± 12.37 72.04 ± 11.25 $<$0.001
Female 0 40 (46.5%) 721 (54.0%) 0.2149
Race 0 $<$0.001
White 70 (81.4%) 1242 (93.0%)
African American 12 (14.0%) 60 (4.5%)
Others 4 (4.6%) 34 (2.5%)
BMI 4 28.40 ± 5.32 26.40 ± 6.20 $<$0.001
Comorbidities
Hypertension 0 85 (99%) 1,323 (99%) 0.863
Diabetes 0 63 (73.3%) 788 (59.0%) 0.009
Anemia 0 55 (64.0%) 828 (62.0%) 0.714
Metabolic acidosis 0 22 (25.6%) 240 (18.0%) 0.077
Proteinuria 0 11 (12.8%) 227 (17.0%) 0.312
Secondary hyperparathyroidism 0 28 (32.6%) 240 (18.0%) $<$0.001
Phosphatemia 0 4 (4.7%) 40 (3.0%) 0.39
Heart failure 0 6 (7.0%) 120 (9.0%) 0.526
Stroke 0 1 (1.2%) 40 (3.0%) 0.506
Conduction & dysrhythmias 0 4 (4.7%) 214 (16.0%) 0.005
Claims-driven features
Count of pharmacy claims 0 120 ± 94 109 ± 86 0.293
Count of inpatient claims 0 3.85 ± 3.41 3.74 ± 3.62 0.773
Count of outpatient claims 0 27.78 ± 24.75 22.07 ± 19.13 0.039
Count of professional claims 0 105.37 ± 77.56 87.43 ± 68.02 0.039
Net cost of pharmacy claims 0 12053 ± 11596 10440 ± 20662 0.242
Net cost of inpatient claims 0 33909 ± 53540 29440 ± 32541 0.446
Net cost of outpatient claims 0 9354 ± 17522 8554 ± 17492 0.682
Net cost of professional claims 0 15512 ± 18657 11640 ± 12748 0.061
Range of claims costs 0 11352 ± 32606 8852 ± 11550 0.481
Standard deviation of claims costs 0 831 ± 1263 757 ± 806 0.593
Clinical-driven features
eGFR 0 17.21 ± 5.46 22.78 ± 5.66 $<$0.001
Hemoglobin 3 12.15 ± 2.19 14.25 ± 1.8 $<$0.001
Bicarbonate 9 22.9 ± 6.36 25.3 ± 4.22 0.001
Serum calcium 6 9.39 ± 3.62 10.21 ± 2.86 0.042
Phosphorus 13 3.61 ± 0.87 3.52 ± 0.72 0.350
CKD stage 3 duration 0 3.7 ± 0.6 3.9 ± 1.4 0.0009
Occurrence of CKD stage 4 0 47 (54.7%) 298 (22.3%) $<$0.001
Occurrence of CKD stage 5 0 42 (48.8%) 277 (20.7%) $<$0.001
Number of emergency department visits 16 2.63 ± 2.18 2.01 ± 1.99 0.005
Table 1 summarizes the demographic, comorbidity, claims-driven, and clinical-driven characteristics of the patient cohort, comprising 1,422 patients; 86 progressed to ESRD (6%), 1,336 did not. For patients who progressed to ESRD, the mean time from CKD stage 3 to ESRD was 4.82 ± 1.82 years. The distribution plot of progression times is provided in Supplementary Figure 1. Patients who progressed to ESRD were significantly younger ( 69.13 vs. 72.04 years, $p<0.001$ ), with notable racial disparities. African American patients were more likely to progress to ESRD (14.0% vs. 4.5%), while White patients had lower progression rates ( $81.4\%$ vs. 93.0%, $p<0.001$ ).
Comorbidities such as hypertension were prevalent in both groups (99%), but diabetes was more frequent in the ESRD group ( $73.3\%$ vs. $0.00959.0\%,p=$ 0.009 ). Secondary hyperparathyroidism and conduction/dysrhythmias were also significantly higher in ESRD patients ( $p<0.001$ and $p=0.005$ , respectively). In claims-driven features, ESRD patients showed slightly higher counts of outpatient and professional claims ( $p=0.039$ ) but no significant differences in costs. Clinical-driven features highlighted lower eGFR (17.21 vs. 22.78, $p<0.001$ ), lower hemoglobin levels (12.15 vs. $0.00114.25,p<$ 0.001 ), and higher occurrences of advanced CKD stages in the ESRD group. These differences underscore the clinical and demographic factors associated with ESRD progression.
5.2 Performance Comparison of Models Across Datasets
To facilitate comprehension of model performance across different data sources, we present results using a 24-month observation window, which yielded optimal performance across most models. Table 2 compares the predictive performance of traditional machine learning and deep learning models across three scenarios: claims data-only, clinical data-only, and merged data.
The results demonstrate several key findings across the different data sources. In the claims data-only scenario, deep learning models, particularly LSTM (AUROC: 0.92, F1: 0.54) and GRU (AUROC: 0.92, F1: 0.50), substantially outperformed traditional machine learning approaches such as logistic regression (AUROC: 0.72, F1: 0.33) and random forest (AUROC: 0.74, F1: 0.36). For clinical data, while the performance gap narrowed, deep learning models maintained their advantage, with LSTM achieving the highest performance (AUROC: 0.88, F1: 0.60). Most notably, the merged dataset combining both claims and clinical data yielded the best overall performance, with LSTM achieving better results across all metrics (AUROC: 0.93, F1: 0.65, AUPRC: 0.61). Comprehensive results across all models and observation windows are provided in Supplementary Table 1.
Table 2: Model performance for prediction of ESRD, using claims data-only, clinical data-only, and merged data.
(a) Model performance for prediction of ESRD by using claims data.
Model Model performance metric
F1 Score AUROC AUPRC
Logistic Regression 0.33 0.72 0.44
Random Forest 0.36 0.74 0.48
XGBoost 0.39 0.75 0.47
CNN 0.45 0.82 0.50
RNN 0.50 0.90 0.52
LSTM 0.54 0.92 0.55
GRU 0.50 0.92 0.53
TCN 0.52 0.88 0.53
(b) Model performance for prediction of ESRD by using clinical data.
Model Model performance metric
F1 Score AUROC AUPRC
Logistic Regression 0.54 0.76 0.47
Random Forest 0.58 0.79 0.51
XGBoost 0.57 0.80 0.52
CNN 0.56 0.84 0.53
RNN 0.61 0.85 0.53
LSTM 0.60 0.88 0.56
GRU 0.60 0.87 0.55
TCN 0.61 0.83 0.54
(c) Model performance for prediction of ESRD by using merged data.
Model Model performance metric
F1 Score AUROC AUPRC
Logistic Regression 0.55 0.75 0.45
Random Forest 0.60 0.84 0.49
XGBoost 0.61 0.85 0.51
CNN 0.56 0.80 0.46
RNN 0.62 0.87 0.53
LSTM 0.65 0.93 0.61
GRU 0.63 0.90 0.58
TCN 0.61 0.89 0.58
5.3 Key Feature Analysis Using Explainable AI Techniques
Having identified the 24-month observation window as optimal and LSTM as the best-performing model for ESRD prediction, we employed two complementary explainable AI techniques to understand the features driving these predictions. At the cohort level, we utilized XGBoost’s feature importance analysis to identify key predictive variables across the population. At the individual level, we applied SHAP analysis to provide detailed, patient-specific insights into model predictions.
5.3.1 Cohort-Level Key Feature Identification
Our feature importance analysis of the optimal ML model, XGBoost, revealed several key predictors for ESRD progression (see Fig.  4(a)). The most critical feature was the presence of stage 5 CKD (S5), aligning with clinical literature and expert feedback on its importance in predicting ESRD. Claims-related features, such as the number of outpatient claims (n_claims_O) and total inpatient claim expenses (net_exp_I), also ranked highly. The diverse types of top-ranked features underscore the benefit of multi-sourced data integration, enhancing the model’s predictive power. This combination of clinical and claims data not only improves the accuracy of the predictions but also supports more comprehensive and reliable decision-making in clinical practice. However, while XGBoost’s feature importance reflects the absolute contribution of each feature, it does not specify whether the influence is positive or negative.
(a) Top 30 features of importance for the XGBoost model.
We present SHAP force plots for two patients, Patient X and Patient Y, both of whom were correctly predicted for their risk of progressing to ESRD. Despite both having CKD Stage 5, they showed notably different patterns of contributing features, highlighting how the same clinical outcome can arise from distinct combinations of risk factors and underscoring the importance of personalized treatment strategies.
This variation in key features, even among patients with the same CKD stage, highlights the importance of aiding the identification of individual differences in treatment planning. SHAP analysis offers a clear visualization of these factors, enabling healthcare providers to tailor interventions to each patient’s unique risk profile and modifiable factors.
5.4 Analysis of Model Misclassifications: Type I and Type II Errors
To explore these misclassifications, we analyzed Type I and Type II errors, as depicted in Figure 5. Notably, most incorrect predictions for patients who progressed to ESRD, yet were predicted otherwise, cluster near the lower end of the plot rather than around the 0.5 decision threshold or randomly scattered. This clustering suggests consistent factors may be influencing these errors, warranting a deeper investigation.
To understand the underlying causes of model misclassification, we conducted a detailed analysis of Type I and Type II errors. We focused particularly on Type II errors, where patients were incorrectly predicted not to develop ESRD but ultimately did progress to the condition. By comparing the 16 false negatives (patients shown in the lower part of the right plot) with the 70 true positives (patients correctly predicted to progress to ESRD, shown in the upper part of the rightplot), we aimed to identify patterns and characteristics that might explain why the model failed to identify these high-risk patients. These errors can delay essential treatment, potentially accelerating disease progression.
Conversely, Type I errors represent patients who were incorrectly predicted to develop ESRD but did not actually progress. Our analysis contrasted 17 false positives (patients shown in the upper part of the left plot) with 1,319 true negatives (patients correctly predicted not to progress, shown in the lower part of the left plot). These Type I errors are particularly concerning in clinical practice as they could lead to unnecessary medical interventions and inefficient allocation of healthcare resources, potentially diverting attention from patients at genuine risk of ESRD progression. Understanding the characteristics distinguishing these misclassified cases from correctly predicted ones may provide valuable insights for model refinement and improved clinical decision support.
Figure 5: Analysis of Model Misclassification: Type I and Type II Errors. The left plots illustrate the predicted probability distribution for all patient outcomes, highlighting the model’s performance across the dataset. The right plots focus on the incorrectly predicted cases, distinguishing between Type I errors (false positives) and Type II errors (false negatives).
Model Mean P-value Timestamp
Correct Incorrect
n_claims_I 0.70 0.14 0.048752 5
n_claims_O 5.71 2.67 0.001325 2
net_exp_O 3831.83 830.61 0.002747 0
S5 0.36 0.06 0.001189 6
S5 0.37 0.49 0.662073 7
net_exp_O 4219.01 1044.37 0.001883 4
Model Mean P-value Timestamp
Correct Incorrect
n_claims_DR 13.52 7.21 0.001156 3
n_claims_O 3.03 1.69 0.009384 4
S4 3831.83 830.61 0.002747 4
S5 0.09 0.35 0.000250 6
S5 0.08 0.35 0.001388 7
net_exp_O 1044.82 333.61 0.000324 4
Based on Table 3(a), CKD Stage 5 emerges as a critical feature. For patients incorrectly predicted as not progressing to ESRD (false negatives), the mean value for CKD Stage 5 at timestamp 6 is 0.06, which is significantly lower than the 0.36 mean for those correctly predicted to progress. This suggests that these patients did not have a Stage 5 record until the subsequent timestamp. The model likely interpreted the absence of a Stage 5 indicator up to timestamp 7 as a sign that ESRD development was unlikely. However, after consulting with nephrologists and examining the literature on CKD progression, we note that some CKD patients may experience a sudden decline in kidney function, ”crashing” into dialysis, where they appear stable for a period before rapidly progressing to ESRD due to acute kidney injury or treatments during hospitalization.
5.5 Impact of Updated eGFR Equation on Racial Bias in Predictions
(c) Demographic distribution patterns under the 2021 eGFR equation. Encoding mappings are: Race: 0=White, 1=Black, 2=Others; Gender: 0=Male, 1=Female; ESRD Status: 0=non-ESRD, 1=ESRD.
When comparing predictions between the 2009 and 2021 eGFR equations (Fig  6(a),  6(b)), we observed a decrease in total incorrect predictions from 33 to 28, reducing the overall error rate from 0.0232 to 0.0228. Among these misclassifications, only 7 cases were common to both equations. Analysis of racial distribution in prediction errors (Fig  6(c),  6(d)) revealed varying patterns across different groups: incorrect predictions for white patients increased from 2 to 4, African American patients decreased from 4 to 1, and patients in the ”Other” category increased from 10 to 12. The demographic breakdown for the 2021 equation cohort, including race, gender, and ESRD outcomes, is shown in Fig  6(c). However, these numbers are clearly too small for robust insights.
6 Discussion
This study demonstrates the value of integrating multiple data sources and deep learning methodologies for improving ESRD progression predictions in CKD patients. By combining claims and EHR data with LSTM and GRU models, which excel at capturing temporal dependencies, we achieved enhanced predictive accuracy using a 24-month observation window. This window length provides an optimal balance between early detection and prediction reliability, supporting timely clinical interventions. Our research addresses a critical gap in chronic disease management by integrating clinical and claims data, offering a more comprehensive view of patient health trajectories than single-source approaches. The combination of medical indicators and healthcare utilization patterns provides deeper insights into ESRD progression risk factors, making the model more robust across diverse patient populations. The determination of this optimal observation window represents a significant advancement in balancing timely interventions with prediction accuracy, enabling healthcare providers to efficiently monitor CKD progression while minimizing data collection burden.
Our findings suggest that while the updated 2021 eGFR equation shows modest improvement in overall prediction accuracy compared to the 2009 equation (error rate reduction from 0.0232 to 0.0228) in this study population, the more useful insight may come from the pattern of misclassifications. The fact that only 7 cases were misclassified by both equations, despite similar overall error rates, indicates that these equations capture different aspects of kidney function and disease progression. This has important implications for clinical practice: while the 2021 equation may better serve minoritized populations by removing race-based adjustments, the persistence of prediction errors clustered in certain patient profiles suggests that eGFR alone may not be sufficient for accurate ESRD prediction. These findings underscore the need for a multi-faceted approach to risk assessment, combining both traditional clinical markers and novel predictive features. By leveraging integrated data sources and advanced modeling techniques, healthcare providers can develop more comprehensive risk assessment strategies, potentially leading to more targeted interventions and improved patient outcomes across diverse populations.
6.1 Limitations
Our study’s reliance on data from a single institution limits the model’s generalizability to all care delivery settings. Incorporating data from more sources could enhance robustness and applicability across different clinical settings. The framework is also currently tailored to predict ESRD outcomes specifically in CKD patients; expanding it to other chronic diseases could provide broader insights.
The use of EHR data presents challenges, such as observational bias, incomplete data, and underrepresentation of certain populations, which can impact model accuracy and fairness. Addressing these issues through comprehensive datasets or advanced imputation techniques is necessary for further improving prediction reliability.
Despite employing data oversampling methods, the disparity between our model’s high AUROC scores and relatively lower F1 and AUPRC scores indicates persistent challenges in addressing data imbalance. This gap suggests that traditional oversampling techniques may not fully compensate for the inherent imbalance in ESRD progression data. To enhance performance, exploring advanced data balancing methods is essential. Techniques such as ensemble methods, cost-sensitive learning, and hybrid approaches that combine oversampling and undersampling have shown promise in effectively managing imbalanced datasets.
6.2 Future Directions
Expansion to Other Chronic Diseases: To evaluate framework’s versatility, we plan to apply it to other chronic conditions, such as heart disease. This will help assess its applicability across various care delivery settings, providing insights into its broader potential for chronic disease management and informing further optimization efforts.
7 Conclusion
8 Acknowledgments
The authors sincerely thank the community nephrology practice for providing access to their de-identified clinical data and the wealth of medical knowledge and inputs shared by several nephrologists in the practice and their CIO on the raw data and the results of our analysis. Similarly, we are very grateful to the health insurance organization for providing access to the de-identified claims data on their shared set of patients with the nephrology practice and for the invaluable help provided by their data team in understanding the claims data.
AI Writing Assistance: All drafts of this manuscript were created by the authors without AI assistance. GPT-4o and Claude 3.5 (Sonnet) were used solely for polishing grammar and refining expression. The outputs from these tools were critically compared and reviewed by the authors to ensure accuracy and clarity. Final revisions were made by the authors to ensure all expressions align with the intended meaning and academic standards.
Funding Statement
This work was supported by the fellowship support for Y.L. from the Center for Machine Learning and Health (CMLH) at Carnegie Mellon University.
Competing Interests Statement
The authors have no competing interests to declare.
Data Availability Statement
The data underlying this article consist of active healthcare clinical and claims data, which cannot be shared publicly due to privacy regulations and the sensitive nature of the information.
Supplementary
Supplementary Table 1: Model Performance Metrics Across Observation Windows
@@ 8 Acknowledgments
Table 4: Comparison of AUROC and F1 scores across different models, using claims data-only, clinical data-only, and merged data. Bold font is used to mark the best performance within each category (ML or DL methods), while bold and italic font highlights the best overall performance across all methods (ML + DL).